# Circumstellar Interaction in the Ultraviolet Spectra of SN 2023ixf 14–66 Days After Explosion

### v1

@@ 5 Comparison with Other Supernovae
Bostroem et al. (2023c) noted that SN 1999em has an extremely low UV flux relative to comparable sequences. It is interesting to consider both the anomalous SED and features of SN 1999em relative to the other supernovae in this sample, as SN 1999em is a relatively prototypical Type IIP supernova (although a high-velocity feature was detected in both $\mathrm{H\alpha}$ and $\mathrm{H\beta}$ during the photospheric phase, interpreted as a sign of CSM interaction; Chugai et al., 2007). The explosion epoch used is half way between the last nondetection and first detection. It is possible that, at discovery, SN 1999em was up to 5 days older, as suggested by Dessart & Hillier (2006), which would make the UV flux less unusual; however, even at a later phase the shape of the UV flux is different from that of the other supernovae — either more suppressed below 2750 Å or with increasing flux redward of that. The presence of an“emission” feature owing to the window of absorption at 1924 Å further supports a later explosion epoch. Given this uncertainty, SN 1999em is plotted in both Figure 6 and Figure 7.
@@
At day 20, the UV flux has decreased in all supernovae. As was observed in SN 2022wsp (Vasylyev et al., 2023b), the Fe feature blueward of Mg II has faded in SN 2023ixf. Interestingly, SN 2021yja shows an extended feature that encompasses both the Fe and Mg II features, possibly indicating that these are blended. The slope and features of SN 2021yja and SN 2022wsp are very similar to the model with $L_{\rm sh}=2.5\times 10^{42}$ $\mathrm{erg~{}s^{-1}}$, although they all have a larger flux. The slope of SN 2023ixf and SN 1999em is not reproduced by any model, even if an overall flux offset is applied. It is possible that models customized for each supernova with more sophisticated physics would provide a better fit. For example, a compelling time-dependent model for SN 1999em is presented in Dessart & Hillier (2006). The flux of SN 2021yja has fallen to a level similar to that of SN 2022wsp, indicating a lower level of power from CSM interaction. The fact that the flux of SN 2021yja fell relative to SN 2023ixf and SN 2022wsp from the flux levels at day 10 gives us confidence that this is not a distance effect. The window of absorption at 1924 Å creates a clear “emission line” in SN 2021yja, SN 2022wsp, and SN 2023ixf.
@@
In interacting supernovae, the Mg II emission is attributed to emission from the cool dense shell which has been excited by radiation from the reverse shock (Chevalier & Fransson, 1994; Dessart & Hillier, 2022). Originating from a shell of material, the emission-line profile is expected to be broad, boxy, and blue shifted (Dessart & Hillier, 2022). It has been noted in SN 1979C 14 yr after explosion (Fesen et al., 1999), in SN 1993J on day 670 (Fransson et al., 2005), in SN 1995N on day 943 (Fransson et al., 2002), in SN 1998S at days 28–485 (Fransson et al., 2005), and in SN 2010jl at days 34–543 (Fransson et al., 2014). However, these supernovae represent the more strongly interacting objects in our sample. It remains to be seen if this feature is present in supernovae which do not have any period of strong interaction with dense CSM.
@@
We searched all available NUV spectra of our comparison sample of Type II supernovae for the broad, boxy Mg II emission that is observed on day 66 in SN 2023ixf. As Mg II is present in absorption at day 25 and then in emission at day 66, we do not know the exact timing of this transition. Additionally, this timing is related to the ejecta and CSM properties and thus may not be a universal property of Type II supernovae. We therefore consider all spectra with phases greater than 30 days. One complicating factor is the presence of Mg II absorption from both the Milky Way and the host galaxy, which contaminates the supernova emission and can make it challenging to assess the shape of the emission profile. Whenever possible, we remove this narrow absorption from our data.
@@
We do not detect this feature in SN 1987A, although this is perhaps not surprising given its blue supergiant progenitor and lack of UV flux. The feature is clearly detected in the remaining supernovae as a broad, boxy, blue-shifted profile. In SN 1980K, we marginally detect it. However, the S/N of the spectra is fairly low and the feature is not as strong as in other supernovae. In SN 1979C we detect it developing already at day 24, although it could be as old as 32 days if the supernova explosion occurred immediately after the last nondetection. From this date, Mg II evolves from a blue-shifted, flat-topped profile into one with slanted profile (higher blue edge) starting around day 60. Similarly, Mg II evolves in SN 1998S from a boxy, blue-shifted, symmetric profile at day 72 to an asymmetric profile at days 237 and 485. In SN 1993J, the Mg II emission is fairly constant with a boxy, asymmetric profile from day 173 to 669. This is the earliest epoch of data for SN 1993J after the UV continuum has faded. While the feature is clearly present as broad emission in SN 1995N and SN 2010jl (days 30–572), the ISM absorption from both the Milky Way and the host galaxy obscure the shape of the profile and we do not include these in our analysis.
@@
The Mg II emission features of SN 1979C, SN 1980K, SN 1993J, and SN 1998S are compared to that of SN 2023ixf in Figure 8. We choose the closest available epoch to day 66, and for SN 1998S we also include the first epoch at which the profile is asymmetric. From these comparisons we find that if a change in the shape of the Mg II emission is observed, the emission begins as a symmetric, boxy, blue-shifted profile and always evolves to the same asymmetric, boxy profile, with more flux on the blue side. For supernovae that, like SN 2023ixf, always show an asymmetric profile, the asymmetry is always higher on the blue side. The consistency of this asymmetry tells us that this is not an effect of asymmetric CSM which we would expect to observe from a different viewing angle, but rather an optical-depth effect. This is corroborated by the Dessart & Hillier (2022) models which show a blue-shifted, boxy emission profile with no emission from ejecta at velocities below 7000 $\mathrm{km~{}s^{-1}}$ (i.e. emission is isolated to the cool dense shell). Additional validation is derived from a custom model reproduces the models of Dessart & Hillier (2022) but with the cool dense shell at 8500 $\mathrm{km~{}s^{-1}}$ and shows the same slanted profile as we observed in the Mg II feature. While optical depth effects are responsible for the blue-shifted and slated profile at early times, at later times dust formation likely attenuates the red side of the feature. This sample shows a variety in maximum velocity of the blue side of these features, indicating diversity in the location of the cool dense shell, as well as other supernova parameters such as ejecta mass, hydrogen envelope mass, explosion energy, and CSM mass. However, they uniformly rise to maximum more slowly than the models, indicating a different cool dense shell density profile than is implemented in the models.
@@
Figure 8: A comparison of the Mg II emission line in SN 2023ixf (day 66; black) to that of other supernovae which show this feature (SN 1979C, brown; SN 1993J, yellow; and SN 1998S, magenta). All lines have a blueshifted peak. In all but the first SN 1998S spectrum, the profiles are asymmetric indicating that the red side of the line is attenuated. The gap in some spectra is due to the removal of narrow Galactic and host ISM absorption.
@@ 6.1 Testing Model Predictions
    A sharp-edged red shelf extending beyond the photospheric emission in $\mathrm{H\alpha}$ originating from the cool dense shell, although this feature is only starting to become visible at day 50
@@
For the first point, we find that the UV luminosity is consistent with models with CSM interaction at all epochs, and not just limited to the early epochs when the narrow emission features are present. The fact that the shock power in the best-matched model decreases with lookback time indicates that the ejecta are sweeping through CSM created from decreasing mass-loss rates. We detect a boxy Mg II emission feature at 66 days, indicating CSM interaction even without the presence of UV continuum, confirming the second expectation. As described in the third point, the absorption components in the P Cygni profiles of the observed spectra are very shallow and slow to develop, especially in $\mathrm{H\alpha}$ which is only clearly identified in the final epoch. The fourth prediction of a sharp red shelf in $\mathrm{H\alpha}$ is subtle at day 66 and not visible prior to this date. At the S/N of our data we cannot confidently identify it nor rule out its presence. Similarly, the narrow high-velocity absorption component, if present, is significantly weaker than predicted in the models. This could be due to a lower density CSM, asymmetric CSM where the bulk of the CSM is offset from the line of sight, and/or the fragmenting of the cool dense shell due to Rayleigh Taylor instabilities which would spread this feature out in velocity space.
@@
To explore whether clumping and lowering the velocity of the cool dense shell would provide a better fit to our observations, we calculate a custom CMFGEN model. We place the cool dense shell at 8500 $\mathrm{km~{}s^{-1}}$. At this velocity, the ejecta are optically thick. This has two effects: first, the material has not been able to cool, so the temperature is higher; and second, it responds much more quickly to the power from the CSM interaction. However, we found that this change alone was not enough to reproduce our observations and continued to further tune the model. To better approximate the physical conditions that create the emission from the cool dense shell (i.e., the reprocessing of X-rays from the forward and reverse shock), we add a uniform field of X-rays starting at 8000 $\mathrm{km~{}s^{-1}}$ with a total power of $5\times 10^{42}$  $\mathrm{erg~{}s^{-1}}$ instead of inserting power into the cool dense shell directly. Given the energy of these X-rays, we add higher energy ions to our model, namely C III, N III, O III, Mg III, and Ne III. We also increase the clumping from 1% to 10% volume filling factor outside of 8000 $\mathrm{km~{}s^{-1}}$, keeping the total mass the same.
@@
In Figure 9 we compare this model at day 14.6 to our day 14.2 observed spectrum and the best-matching model from Dessart & Hillier (2022) at day 14.6. Reducing the velocity of the cool dense shell aligns the absorption and emission. However, the resulting increase in the temperature of the cool dense shell shifts flux from the optical to the UV, causing the UV to overshoot the observed spectrum and the optical to underestimate the true flux. The change in power deposition from direct injection into the cool dense shell to an X-ray field reduces the UV flux to align with the observations but does not resolve the low optical flux issue. The increase in clumping from a volume filling factor of 1% to 10% increases the optical flux to match the full SED. The observed spectrum is very well matched by our custom model, demonstrating the success of this approach. However, further improvements could be made to better match the strength of features in the UV, most notably the Fe II absorption complex in the range 2250–2760 Å which is too strong in the models.
@@
Figure 9: The observed spectrum at day 14 compared to the $L_{sh}=5\times 10^{42}$  $\mathrm{erg~{}s^{-1}}$ model from Dessart & Hillier (2022) (blue) and our new model with the cool dense shell at lower velocity, increased clumping, and a more complete atomic model (pink). The new model is able to match the velocity of most of the features and provides a significantly better match to the SED.
@@ 6.2 CSM Density Profile and Progenitor Mass-loss History
Using the best models at each epoch, we derive the density profile of the RSG progenitor of SN 2023ixf, which is given in Table 2 and plotted in Figure 10. We use the shock luminosity of each model, assuming a shock velocity of $v_{\rm sh}=10,000$ $\mathrm{km~{}s^{-1}}$ to find the CSM density at each epoch from $\rho(r)=L_{\rm sh}/(2\pi v_{\rm sh}^{3}r^{2})~{}\mathrm{g~{}cm^{-3}}$. Note that this is a simplified model which assumes both spherical symmetry and a shock velocity. From this we find that the density decreases more rapidly than a constant wind, going from $\rho=5.2\times 10^{-16}$ $\mathrm{g~{}cm^{-3}}$ at 14 days after explosion to $\rho=4.8\times 10^{-20}$ $\mathrm{g~{}cm^{-3}}$ at day 66.
@@
If we further use the RSG wind velocity of $v_{\rm wind}=55$ $\mathrm{km~{}s^{-1}}$ measured by Zhang et al. (2023), we can derive a mass-loss rate from the density and a time of mass loss from the radius (see Figure 11). We find $\dot{M}=8.7\times 10^{-4}$ $\mathrm{M_{\odot}~{}yr^{-1}}$ at 14 days after explosion, decreasing to $4.4\times 10^{-4}$ $\mathrm{M_{\odot}~{}yr^{-1}}$ 19 and 24 days after explosion, and finally arriving at a canonical RSG wind of $1.7\times 10^{-6}$ $\mathrm{M_{\odot}~{}yr^{-1}}$ at 66 days after explosion. This traces the mass loss from elevated rates $\sim 7$ yr before explosion down to nominal RSG rates $\sim 35$ yr before explosion. The clear presence of Mg II emission at day 66 indicates that we can detect ejecta-CSM interaction, even with canonical RSG winds of $10^{-6}$ $\mathrm{M_{\odot}~{}yr^{-1}}$. This mass-loss rate and timing is dependent on the assumed wind velocity. If Zhang et al. (2023) measured a wind that had already been radiatively accelerated, the velocity could be lower, which would increase the period of time prior to explosion probed by these observations and decrease the mass-loss rate.
@@
Figure 10: The CSM density of SN 2023ixf as a function of radius assuming a shock velocity of 10,000 $\mathrm{km~{}s^{-1}}$. The CSM density constraints from this paper are plotted with star symbols. The filled symbols are constrained by the overall UV and optical flux. The open stars represent the value from the Mg II line which lies between $L_{\rm sh}=10^{40}$  $\mathrm{erg~{}s^{-1}}$ (black) and $L_{\rm sh}=10^{41}$  $\mathrm{erg~{}s^{-1}}$ (pink). Density constraints from flash spectroscopy are shown at the epoch of the first spectrum, day 4, when the majority of the narrow features disappear, and day 7, when all of the narrow features disappear (Bostroem et al., 2023a; Jacobson-Galán et al., 2023; Zhang et al., 2023). Mustard triangles represent a mass-loss rate of $10^{-2}$  $\mathrm{M_{\odot}~{}yr^{-1}}$ and green inverted triangles represent a mass-loss rate of $10^{-3}$  $\mathrm{M_{\odot}~{}yr^{-1}}$. Values derived from the $\mathrm{H\alpha}$ luminosity from Zhang et al. (2023) are shown as light-blue solid squares and the value from Bostroem et al. (2023a) is a light-blue open square. CSM density from NuSTAR X-ray observations (Grefenstette et al., 2023) are plotted as filled blue circles and X-ray observations from Chandra (Chandra et al., 2024) are shown as open blue circles. Lines of constant density corresponding to mass-loss rates of $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, $10^{-5}$, $10^{-6}$ $\mathrm{M_{\odot}~{}yr^{-1}}$ are shown as solid gray lines. While there is significant scatter in the density, it is clear that density is decreasing more rapidly than would be expected from constant mass-loss rates down to the density expected for canonical RSG winds.
@@
These combined measurements show that the mass-loss rate of the progenitor of SN 2023ixf was relatively low until quite close to the time of core collapse ($R\lesssim 5\times 10^{15}$ cm or $t\lesssim 33$ yr prior to explosion), at which point it increased to the large densities inferred from flash-feature observations. However, this figure also highlights that the systematics, uncertainties, and simplifying assumptions in these different techniques lead to one order of magnitude uncertainties in mass-loss rates and densities.
Figure 11: The same as Figure 10 for the mass-loss history of SN 2023ixf assuming a shock velocity ($v_{\rm sh}=10,000$ $\mathrm{km~{}s^{-1}}$) and the CSM wind velocity ($v_{\rm wind}=55$ $\mathrm{km~{}s^{-1}}$). In its last $\sim$30 years of life, the mass-loss rate increased from a nominal RSG wind of $\sim 10^{-6}$ $\mathrm{M_{\odot}~{}yr^{-1}}$ to $\sim 10^{-2}$ $\mathrm{M_{\odot}~{}yr^{-1}}$.

### v2

@@ 5 Comparison with Other Supernovae
Bostroem et al. (2023c) noted that SN 1999em has an extremely low UV flux relative to comparable sequences. It is interesting to consider both the anomalous SED and features of SN 1999em relative to the other supernovae in this sample, as SN 1999em is a relatively prototypical Type IIP supernova (although a high-velocity feature was detected in both $\mathrm{H\alpha}$ and $\mathrm{H\beta}$ during the photospheric phase, interpreted as a sign of CSM interaction; Chugai et al., 2007). The explosion epoch used is halfway between the last nondetection and first detection. It is possible that, at discovery, SN 1999em was up to 5 days older, as suggested by Dessart & Hillier (2006), which would make the UV flux less unusual; however, even at a later phase, the shape of the UV flux is different from that of the other supernovae — either more suppressed below 2750 Å or with increasing flux redward of that. The presence of an“emission” feature owing to the window of absorption at 1924 Å further supports a later explosion epoch. Given this uncertainty, SN 1999em is plotted in both Figure 6 and Figure 7.
@@
On day 20, the UV flux has decreased in all supernovae. As was observed in SN 2022wsp (Vasylyev et al., 2023a), the Fe feature blueward of Mg II has faded in SN 2023ixf. Interestingly, SN 2021yja shows an extended feature that encompasses both the Fe and Mg II features, possibly indicating that these are blended. The slope and features of SN 2021yja and SN 2022wsp are very similar to the model with $L_{\rm sh}=2.5\times 10^{42}$ $\mathrm{erg~{}s^{-1}}$, although they all have a larger flux. The slope of SN 2023ixf and SN 1999em is not reproduced by any model, even if an overall flux offset is applied. It is possible that models customized for each supernova with more sophisticated physics would provide a better fit. For example, a compelling time-dependent model for SN 1999em is presented in Dessart & Hillier (2006). The flux of SN 2021yja has fallen to a level similar to that of SN 2022wsp, indicating a lower level of power from CSM interaction. The fact that the flux of SN 2021yja fell relative to SN 2023ixf and SN 2022wsp from the flux levels on day 10 gives us confidence that this is not a distance effect. The window of absorption at 1924 Å creates a clear “emission line” in SN 2021yja, SN 2022wsp, and SN 2023ixf.
@@
In interacting supernovae, the Mg II emission is attributed to emission from the cool dense shell, which has been excited by radiation from the reverse shock (Chevalier & Fransson, 1994; Dessart & Hillier, 2022). Originating from a shell of material, the emission-line profile is expected to be broad, boxy, and blue shifted (Dessart & Hillier, 2022). It has been noted in SN 1979C 14 yr after explosion (Fesen et al., 1999), in SN 1993J on day 670 (Fransson et al., 2005), in SN 1995N on day 943 (Fransson et al., 2002), in SN 1998S on days 28–485 (Fransson et al., 2005), and in SN 2010jl on days 34–543 (Fransson et al., 2014). However, these supernovae represent the more strongly interacting objects in our sample. It remains to be seen if this feature is present in supernovae that do not have any period of strong interaction with dense CSM.
@@
We searched all available NUV spectra of our comparison sample of Type II supernovae for the broad, boxy Mg II emission that was observed on day 66 in SN 2023ixf. As Mg II is present in absorption on day 25 and then in emission on day 66, we do not know the exact timing of this transition. Additionally, this timing is related to the ejecta and CSM properties and thus may not be a universal property of Type II supernovae. We therefore consider all spectra with phases greater than 30 days. One complicating factor is the presence of Mg II absorption from both the MW and the host galaxy, which contaminates the supernova emission and can make it challenging to assess the shape of the emission profile. Whenever possible, we remove this narrow absorption from our data.
@@
We do not detect this feature in SN 1987A, although this is perhaps not surprising given its blue supergiant progenitor and lack of UV flux. The feature is clearly detected in the remaining supernovae as a broad, boxy, blue-shifted profile. In SN 1980K, we marginally detect it. However, the S/N of the spectra is fairly low and the feature is not as strong as in other supernovae. In SN 1979C we detect it developing already on day 24, although it could be as old as 32 days if the supernova explosion occurred immediately after the last nondetection. From this date, Mg II evolves from a blue-shifted, flat-topped profile into one with a slanted profile (higher blue edge) starting around day 60. Similarly, Mg II evolves in SN 1998S from a boxy, blue-shifted, symmetric profile on day 72 to an asymmetric profile on days 237 and 485. In SN 1993J, the Mg II emission is fairly constant with a boxy, asymmetric profile from day 173 to 669. This is the earliest epoch of data for SN 1993J after the UV continuum has faded. While the feature is clearly present as broad emission in SN 1995N and SN 2010jl (days 30–572), the ISM absorption from both the Milky Way and the host galaxy obscure the shape of the profile and we do not include these in our analysis.
@@
The Mg II emission features of SN 1979C, SN 1980K, SN 1993J, and SN 1998S are compared to that of SN 2023ixf in Figure 8. We choose the closest available epoch to day 66, and for SN 1998S we also include the first epoch at which the profile is asymmetric. From these comparisons, we find that if a change in the shape of the Mg II emission is observed, the emission begins as a symmetric, boxy, blue-shifted profile and always evolves to the same asymmetric, boxy profile, with more flux on the blue side. For supernovae that, like SN 2023ixf, always show an asymmetric profile, the asymmetry is always higher on the blue side. The consistency of this asymmetry tells us that this is not an effect of asymmetric CSM that we would expect to observe from a different viewing angle, but rather an optical-depth effect. This is corroborated by the Dessart & Hillier (2022) models, which show a blue-shifted, boxy emission profile with no emission from ejecta at velocities below 7000 $\mathrm{km~{}s^{-1}}$ (i.e. the emission is isolated to the cool dense shell). Additional validation is derived from a custom model that reproduces the models of Dessart & Hillier (2022) but with the cool dense shell at 8500 $\mathrm{km~{}s^{-1}}$ and shows the same slanted profile as we observed in the Mg II feature. While optical-depth effects are responsible for the blue-shifted and slated profile at early times, at later times dust formation likely attenuates the red side of the feature. This sample shows a variety in maximum velocity of the blue side of these features, indicating diversity in the location of the cool dense shell, as well as other supernova parameters such as the ejecta mass, hydrogen envelope mass, explosion energy, and CSM mass. However, they uniformly rise to maximum more slowly than the models, indicating a different cool dense shell density profile than is implemented in the models.
@@
Figure 8: Comparison of the Mg II emission line in SN 2023ixf (day 66; black) to that of other supernovae that show this feature (SN 1979C, brown; SN 1993J, yellow; and SN 1998S, magenta). All lines have a blueshifted peak. In all but the first SN 1998S spectrum, the profiles are asymmetric indicating that the red side of the line is attenuated. The gap in some spectra is due to the removal of narrow Galactic and host ISM absorption.
@@ 6.1 Testing Model Predictions
    A sharp-edged red shelf extending beyond the photospheric emission in $\mathrm{H\alpha}$ originating from the cool dense shell, although this feature is only starting to become visible on day 50
@@
For the first point, we find that the UV luminosity is consistent with models with CSM interaction at all epochs, and not just limited to the early epochs when the narrow emission features are present. The fact that the shock power in the best-matched model decreases with the lookback time indicates that the ejecta are sweeping through CSM created from decreasing mass-loss rates. This is consistent with the light-curve modeling results of Singh et al. (2024) which require extended CSM ($\mathrm{\dot{M}=10^{-4}}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$to $\mathrm{R=10^{16}}$ cm) to match the late-time UV excess. We detect a boxy Mg II emission feature at 66 days, indicating CSM interaction even without the presence of a UV continuum, confirming the second expectation. As described in the third point, the absorption components in the P Cygni profiles of the observed spectra are very shallow and slow to develop, especially in $\mathrm{H\alpha}$ which is only clearly identified in the final epoch. This is also notes by Singh et al. (2024). The fourth prediction of a sharp red shelf in $\mathrm{H\alpha}$ is subtle on day 66 and not visible prior to this date. At the S/N of our data, we cannot confidently identify it nor rule out its presence. Similarly, the narrow high-velocity absorption component, if present, is significantly weaker than predicted in the models. This could be due to a lower density CSM, asymmetric CSM where the bulk of the CSM is offset from the line of sight, and/or the fragmenting of the cool dense shell due to Rayleigh Taylor instabilities which would spread this feature out in the velocity space.
@@
To explore whether clumping and lowering the velocity of the cool dense shell would provide a better fit to our observations, we calculate a custom CMFGEN model. We place the cool dense shell at 8500 $\mathrm{km~{}s^{-1}}$. At this velocity, the ejecta are optically thick. This has two effects: first, the material has not been able to cool, so the temperature is higher; and second, it responds much more quickly to the power from the CSM interaction. However, we found that this change alone was not enough to reproduce our observations and continued to further tune the model. To better approximate the physical conditions that create the emission from the cool dense shell (i.e., the reprocessing of X-rays from the forward and reverse shock), we add a uniform field of X-rays starting at 8000 $\mathrm{km~{}s^{-1}}$ with a total power of $5\times 10^{42}$  $\mathrm{erg~{}s^{-1}}$ instead of inserting power into the cool dense shell directly. Given the energy of these X-rays, we add higher energy ions to our model, namely, C III, N III, O III, Mg III, and Ne III. We also increase the clumping from 1% to 10% volume filling factor outside of 8000 $\mathrm{km~{}s^{-1}}$, keeping the total mass the same.
@@
In Figure 9 we compare this model on day 14.6 to our day 14.2 observed spectrum and the best-matching model from Dessart & Hillier (2022) on day 14.6. Reducing the velocity of the cool dense shell aligns the absorption and emission. However, the resulting increase in the temperature of the cool dense shell shifts flux from the optical to the UV, causing the UV to overshoot the observed spectrum and the optical to underestimate the true flux. The change in power deposition from direct injection into the cool dense shell to an X-ray field reduces the UV flux to align with the observations but does not resolve the low optical flux issue. The increase in clumping from a volume filling factor of 1%–10% increases the optical flux to match the full SED. The observed spectrum is very well matched by our custom model, demonstrating the success of this approach. However, further improvements could be made to better match the strength of features in the UV, most notably the Fe II absorption complex in the range 2250–2760 Å, which is too strong in the models.
@@
Figure 9: Observed spectrum on day 14 compared to the $L_{\mathrm{sh}}=5\times 10^{42}$  $\mathrm{erg~{}s^{-1}}$ model from Dessart & Hillier (2022, blue) and our new model with the cool dense shell at lower velocity, increased clumping, and a more complete atomic model (pink). The new model is able to match the velocity of most of the features and provides a significantly better match to the SED.
@@ 6.2 CSM Density Profile and Progenitor Mass-loss History
Using the best models at each epoch, we derive the density profile of the RSG progenitor of SN 2023ixf, which is given in Table 2 and plotted in Figure 10. We use the shock luminosity of each model, assuming a shock velocity of $v_{\rm sh}=10,000$ $\mathrm{km~{}s^{-1}}$ to find the CSM density at each epoch from $\rho(r)=L_{\rm sh}/(2\pi v_{\rm sh}^{3}r^{2})~{}\mathrm{g~{}cm^{-3}}$. Note that this is a simplified model that assumes both spherical symmetry and a shock velocity. From this, we find that the density decreases more rapidly than a constant wind, going from $\rho=5.2\times 10^{-16}$ $\mathrm{g~{}cm^{-3}}$ at 14 days after explosion to $\rho=4.8\times 10^{-20}$ $\mathrm{g~{}cm^{-3}}$ on day 66.
@@
If we further use the RSG wind velocity of $v_{\rm wind}=55$ $\mathrm{km~{}s^{-1}}$ measured by Zhang et al. (2023), we can derive a mass-loss rate from the density and a time of mass loss from the radius (see Figure 11). We find $\dot{M}=8.7\times 10^{-4}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$ 14 days after the explosion, decreasing to $4.4\times 10^{-4}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$ 19 and 24 days after the explosion, and finally arriving at a canonical RSG wind of $1.7\times 10^{-6}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$ at 66 days after the explosion. This traces the mass loss from elevated rates $\sim 7$ yr before the explosion down to nominal RSG rates $\sim 35$ yr before explosion. The clear presence of Mg II emission on day 66 indicates that we can detect ejecta-CSM interaction, even with canonical RSG winds of $10^{-6}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$. This mass-loss rate and timing are dependent on the assumed wind velocity. If Zhang et al. (2023) measured a wind that had already been radiatively accelerated, the velocity could be lower, which would increase the period of time prior to the explosion probed by these observations and decrease the mass-loss rate.
@@
Figure 10: CSM density of SN 2023ixf as a function of the radius assuming a shock velocity of 10,000 $\mathrm{km~{}s^{-1}}$. The CSM density constraints from this paper are plotted with star symbols. The filled symbols are constrained by the overall UV and optical flux. The open stars represent the value from the Mg II line, which lies between $L_{\rm sh}=10^{40}$  $\mathrm{erg~{}s^{-1}}$ (black) and $L_{\rm sh}=10^{41}$  $\mathrm{erg~{}s^{-1}}$ (pink). Density constraints from flash spectroscopy are shown at the epoch of the first spectrum, day 4, when the majority of the narrow features disappear, and day 7, when all of the narrow features disappear (Bostroem et al., 2023a; Jacobson-Galán et al., 2023; Zhang et al., 2023). Mustard triangles represent a mass-loss rate of $10^{-2}$  $M\mathrm{{}_{\odot}~{}yr^{-1}}$ and green inverted triangles represent a mass-loss rate of $10^{-3}$  $M\mathrm{{}_{\odot}~{}yr^{-1}}$. Values derived from the $\mathrm{H\alpha}$ luminosity from Zhang et al. (2023) are shown as light-blue solid squares and the value from Bostroem et al. (2023a) is a light-blue open square. CSM density from NuSTAR X-ray observations (Grefenstette et al., 2023) are plotted as filled blue circles and X-ray observations from Chandra (Chandra et al., 2024) are shown as open blue circles. Lines of constant density corresponding to mass-loss rates of $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, $10^{-5}$, and $10^{-6}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$ are shown as solid gray lines. While there is significant scatter in the density, it is clear that density is decreasing more rapidly than would be expected from constant mass-loss rates down to the density expected for canonical RSG winds.
@@
These combined measurements show that the mass-loss rate of the progenitor of SN 2023ixf was relatively low until quite close to the time of core collapse ($R\lesssim 5\times 10^{15}$ cm or $t\lesssim 33$ yr prior to explosion), at which point it increased to the large densities inferred from flash-feature observations. However, this figure also highlights that the systematics, uncertainties, and simplifying assumptions in these different techniques lead to 1 order of magnitude uncertainties in mass-loss rates and densities.
Figure 11: Same as Figure 10 for the mass-loss history of SN 2023ixf assuming a shock velocity ($v_{\rm sh}=10,000$ $\mathrm{km~{}s^{-1}}$) and the CSM wind velocity ($v_{\rm wind}=55$ $\mathrm{km~{}s^{-1}}$). In its last $\sim$30 yr of life, the mass-loss rate increased from a nominal RSG wind of $\sim 10^{-6}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$ to $\sim 10^{-2}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$.